OpenStudy (marissag):

Determine if line l1 [x,y,z]=[4,-3,2]+t[1,8,-3] intersects l2 defined by [x,y,z]=[2,-19,8]+s[4,-5,-9].

OpenStudy (anonymous):

set the components equal to each other. if the s and t you find from the first two equation satisfies the third equation as well, that's the point of intersection.

OpenStudy (marissag):

So, I would end up with 4+t=2+4s, -3+8t=(-19)-5s and 2-3t=8-9t, and then solve from there?

OpenStudy (anonymous):

yep